1. 4 ln x = 5
2. log (x-9) = 1 - log x
have trouble solving this two problems

Answer by stanbon(75887)

1. 4 ln x = 5
ln(x) = 5/4
x = e^(5/4)
x = 3.4903...
----------------------------------
2. log (x-9) = 1 - log x
log(x-9) + log(x) = 1
log[x(x-9)] = 1
log(x^2-9x) = 1
x^2-9x = 10
x^2 - 9x -10 = 0
(x-10)(x+1) = 0
Positive solution:
x = 10
============================
Cheers,
Stan H.